OpenStudy (anonymous):

the product of two consecutive even integers is 288. find the integers??? please help online assignment due!!!

OpenStudy (anonymous):

let these numbers be x and x+2 x(x+2)=288

OpenStudy (anonymous):

x^2+2x=288

OpenStudy (anonymous):

yes, then x^2+2x-288=0 factor

OpenStudy (anonymous):

18, 16 ??

OpenStudy (anonymous):

let these numbers be x and x+2 x(x+2)=288

OpenStudy (anonymous):

right, and also -16, -18, because it wasn't specified that positive even integers were required

OpenStudy (anonymous):

thank you!!!!!